OpenStudy (anonymous):

solove : (x+3)(x-5)(x+4)

OpenStudy (cggurumanjunath):

X= -3,5,-4

OpenStudy (lucaz):

if you want the roots they are the values of x that make this expression equal to zero, {-3,5,-4}

OpenStudy (cggurumanjunath):

(X+3)=0 (X-5)=0 (X+4)=0